OpenStudy (anonymous):

Are the graphs of the lines in the pair parallel? Explain. y = x + 8 –2x + 12y = –11

OpenStudy (anonymous):

A. Yes, since the slopes are the same and the y-intercepts are the same. B. No, since the y-intercepts are different. C. Yes, since the slopes are the same and the y-intercepts are different. D. No, since the slopes are different.

OpenStudy (kropot72):

Rearrange the second equation as follows: 12y = 2x - 11 Then divide by 12: y = x/6 - 11/12 What is the gradient of the last equation?

OpenStudy (anonymous):

y = x + 8 so the slope is 1 -2x + 12y = -11 12y = -11 +2x y = -11/12 + x/6 y = x/6 - 11/12 .... this slope is not equal with the upper one . it cant be parallel so.. D

OpenStudy (anonymous):

b.?

OpenStudy (kropot72):

The gradient of the last equation is 1/6. The gradient of the first equation is 1. So the correct choice is D.

OpenStudy (anonymous):

Yeah. Just like i said be4 :)

OpenStudy (anonymous):

mind helping me with 9 more like this?